### v1

# The 4DLO and other tubing models of $S^{3}$ symmetry

Abstract
The Four-dimensional Light Orchestra or 4DLO was an interactive sculpture at the National Museum of Mathematics (MoMath) from November 20, 2025 through January 2026, illustrating various sub-symmetries of the 24-cell with colored lights, part of a larger sequence of tubing sculptures aiming to bring to life a few lines of tables appearing in [3], reprinted in [4], further illuminated in [11]. Best of all museum patrons could manipulate 4DLO’s lighting by singing and making funny noises into a microphone, and they did with gusto. Here we describe some of the technical aspects of this sculpture and its context.
Figure 1: (left) 4DLO showing a compound of three tesseracts, with student built work
just visible in the back. (right) Singing affected the sculpture and its program.
(a) Refer to caption (b) Refer to caption (c) Refer to caption (d) Refer to caption
Figure 2: (a) The fully symmetries of 24-cell; (b) half those, directing the edges; (c) A transitional moment during a sequence showing off the twenty four octahedral cells of the 24-cell, one by one. A similar sequence showed off the twenty four cubes of the three tesseracts sharing these edges. (d) Four mutually perpendicular groups of four rings of length six.
The 24-cell, also known as an “octaplex” and by other names, is a “four-dimensional regular polyhedron” or regular polytope, described more fully in a moment, denoted $\{3,4,3\}$ by Ludwig Schläfli who enumerated all regular polytopes in all dimensions in 1852. These are akin to the five regular (three-dimensional) polyhedra and to the infinitely many regular polygons in the plane. In each dimension $d\geq 5$ there are disappointingly only three regular polytopes, generalizations of the triangle, tetrahedron, etc., (denoted $\{3,...,3\}$ with one or more 3’s); the square, cube, tesseract, etc., ($\{...,3,4\}$ with no or more 3’s); and the polytopes dual to those, a tilted square, an octahedron, the 16-cell, etc. ($\{4,3,$...\} with no or more 3’s). In four dimensions, for whatever reasons, there are six.
Their enumeration was rediscovered repeatedly; George Hart has a helpful survey (with beautiful 3D printed models and a full bibliography) at [6], H.S.M. Coxeter’s Regular Polytopes [5] is a canonical reference, and the Wikipedia article is extensive with detailed technical information [15]. Schleimer and Segerman’s 2012 Bridges paper [13] gives precise definitions of many terms we use here.
Just as a polyhedron has two-dimensional polygons for faces, and can be considered as a division of the three dimensional sphere, a (four-d) polytope is formed of three-dimensional cells and may be viewed as a division of the hypersphere $S^{3}$. In the 24-cell there are twenty four three-dimensional octahedral cells. Each octahedron (denoted $\{3,4\}$) is composed of triangles $\{3\}$ meeting as though coming together through the sides of a square $\{4\}$. In a similar manner, at each vertex of the 24-cell, there are six octahedral cells, meeting tip to tip, as though they were coming through the faces of a cube $\{4,3\}$. Schläfli’s notation encodes this data as $\{3,4,3\}$.
This shape and further shapes derived from it have been popular, appearing over many Bridges galleries and proceedings, in various forms and to various ends: as collections of monkeys [7], as a beautiful student-created aluminum frame sculpture [2], as balloon art [1], on an engraved pendant [2], and as a basis for more abstract construction [8]. A stainless steel 24-cell memorial radiates in the mathematics department at Pennsylvania State University [9].
Clearly the 24-cell is an appealing shape. Mathematically it is highly symmetrical and an outlier among the regular polytopes, with no clear analogue in other dimensions. Like the tetrahedra-type polytopes of all dimensions, the 24-cell is self-dual — at the center of each of its twenty four cells there is a vertex of another 24-cell, its dual, and vice versa. For the same reason, it has the same number of two dimensional faces as it has edges, 96. Because all these polytopes have palindromic Schläfli symbols, and reversing the symbol of a regular polytope produces its dual, these are all duals of themselves, as are all of the regular polygons, which have symbols of the form $\{n\}$. Two copies of any of these self-dual polytopes form a natural compound. For example, the compound of two self-dual three-dimensional tetrahedra is well-known as Kepler’s Stella Octangula and two dual triangles form the familiar Star of David. So too is there a natural compound of two dual copies of the 24-cell — this structure appears in one of Rinus Roelef’s steel sculptures, landing a photo of the first author and Edmund Harriss in a newspaper [12]. Among these self-dual compounds, only the 24-cell has much structure to its symmetries.
The symmetries of the hypersphere were enumerated in stages, begun by Goursat (1889), continued by Threlfall and Seifert (1931-3), Hurley (1951), Du Val (1964), and Conway and Smith (2003 [3]). In their recent monograph, Rastanawi and Rote give a geometric re-enumeration, which we lean on heavily for evaluating symmetry types [11]. (Their Section 3.3 gives a detailed survey of this history.) In effect, this project is to illustrate part of Conway and Smith’s tables decades late for The Symmetries of Things [4].
Figure 3: (left) A spherical octahedron and its stereographic projection. Both are divisions of their space into triangular cells bounded by circular arcs meeting at right angles. (Right) A stereographic projection of a 16-cell, with tetrahedral cells and faces and edges meeting at right angles.
The problem of visualizing the hypersphere is fundamentally three-dimensional, just as features on the surface of the Earth may be located with just two dimensions (say, latitude and longtitude). Nearly two thousand years ago Claudius Ptolemy produced the first atlas, with thousands of located sites spanning from the West African coast to the South China Sea. He projected points on Earth to a map, that is, from points on the sphere to points in a plane. His conical and cylindrical projections are foundational to cartographers to this day, but they’ve found little use for his stereographic projection despite it being very good for small regions away from some chosen point $N$. It is beloved by mathematicians, however, for not only does it preserves angles as measured on the sphere when they are measured on the plane, but it also takes circles on the sphere to circles in the plane (understanding, as did Nicolaus of Cusa ca. 1450, that lines are simply circles of infinite radius). And this mapping is elegantly defined: Geometrically, we choose some plane $P$ and a point $N$ on a sphere. Stereographic projection takes each point $p$ on the sphere to a point $\pi(p)$ in $P$, defined as the point where the line through $p$ and $N$ meets $P$. ($\pi(N)$ is itself “at infinity,” a kind of NaN.) In Figure 3 this is shown in the standard way, with $N$ at the top and $P$ tangent at N-N. In coordinates, say p=(x,y,,z,w)p=(x,y,...,z,w) of any dimension $d$, it is convenient to take N=(0,..,0,1)N=(0,..,0,-1) and project to the plane with $w=0$, using $\pi(p)=(x,...,z)/(1+w)$, a (d1)(d-1)-dimensional coordinate in $P$.
By examining a pattern in $P$, such as the division of the plane into eight right-angled triangular regions at left in Figure 3, we can understand at least something about a division of the sphere in the same way (though a spherical octahedron will have much more symmetry than any planar drawing of it). We, as have many others, apply the same strategy one dimension up to learn something of the possible symmetries of patterns in the hypersphere.
Figure 4: (left and middle) Two cell-down 24-cells, shown as a compound of three tesseracts and as $T^{*}$ shown as three-fold rotational symmetries of a tetrahedron. [New pix coming] These models [will be] were built at the seventeenth Gathering for Gardner, Feb. 2026, with Scott Vorthmann.
Tubing is manufactured in circular coils, and therefore pieces of it relax into circular arcs — this insight, that tubing is great for stereographic renderings of circles in the hypersphere, was the aha moment that spurred on this entire multifaceted project. With no delay the first author sought what might be at hand, finding some vinyl tubing at an automotive store and six-way connectors as playing jacks at a party store. The first of these models, a 16-cell, is at right in Figure 3, and the second is at left bottom in Figure 5.
[…]
(Right) A Compound of Two Compounds of Three Sixteen Cells, G4G12 (2016).
We describe coordinates as quaternions, indicating $(a,b,c,d)$ as ${\bf{q}}=a{\bf{i}}+b{\bf{j}}+c{\bf{k}}+d{\bf{1}}$. On these we allow the usual rules of arithmetic, with multiplicative identity $\bf{1}$}, plus ${\bf{i}}{\bf{j}}={\bf{k}}$ and the rule that, for any imaginary (that is, ${\text{re}}({\bf{q}}):=d$ is zero) and unit ($a^{2}+b^{2}+c^{2}+d^{2}=1$) quaternion $\bf{q}$}, we have ${\bf{q}}{\bf{q}}=-{\bf{1}}$. (See [13] for further details.) The unit quaternions are exactly the points of the hypersphere $S^{3}$, and therefore coordinatize it. For illustrations, stereographically projecting from $\bf{1}$} into our space places $\bf{1}$} at the origin and $\bf{1}$} “at infinity”, the unit imaginary quaternions on the unit sphere and $\pm{\bf{i}},\pm{\bf{j}}$ and $\pm{\bf{k}}$ upon it as coordinate axes. An active reader might sketch on such a template. The familiar “cell-down” tesseract $\{4,3,3\}$ has sixteen vertices, we’ll call $V_{16}$, at points of the form $\frac{1}{2}(\pm{\bf{i}}\pm{\bf{j}}\pm{\bf{k}}\pm{\bf{1}})$, which in stereographic projection lie at the corners of two nested cubes to either side ($\pm{\bf{1}}$) of the unit sphere. We’ll further sort these points into the “even” ones $V_{16}^{+}$ with an even number of -’s, and its “odd” complement $V_{16}^{-}$. The edges of a tesseract connect vertices that differ by exactly one coordinate, and therefore each edge has one end in $V_{16}^{+}$ and the other in $V_{16}^{-}$. Three tesseracts can be arranged as at left in Figure 1; the yellow one, though tilted to the floor, is cell-down (a cube is at its center) and has vertices in $V_{16}$. It is dual to a “vertex-down” 16-cell with eight vertices, $V_{8}=\{\pm{\bf{1}},\pm{\bf{i}},\pm{\bf{j}},\pm{\bf{k}}\}$. You can quickly sketch a 16-cell of your own by connecting up these points on your template; or see the red 16-cell at far upper left Figure 5. A “cell-down” 16-cell with vertices in $V_{16}^{+}$ is at right in Figure 3 and is lit up at right in Figure 9. In fact $V_{16}\cup V_{8}$, which we denote $V_{24}$, form the vertices a 24-cell, which we’ll call “vertex-down”; its dual “cell-down” 24-cell has vertices $V_{24}^{\prime}$ of the form $\frac{1}{\sqrt{2}}(\pm{\bf{x}}\pm{\bf{y}})$ where ${\bf{x}},{\bf{y}}$ are two of ${\bf{1}},{\bf{i}},{\bf{j}}$, and $\bf{k}$}. The 4DLO of Figures 1 and 2 is vertex-down, but with the coordinate axes tilted in space. The 24-cells of Figure 4, [to be] made at G4G in February 2026, are cell-down.
Figure 6: (left) an obliquely arrangement of three 16-cells; (middle) with directed edges, symmetry $O^{*}$; (right) a pleasing portion of three 16-cells made from four-way and two-way connectors.
[…]
Nicely, quaternionic multiplication can describe symmetries like these: Each given unit quaternion ${\bf{q}}\in S^{3}$ also acts upon on $S^{3}$ as a whole, sliding everything together along a twisting family of great-circle rails (a “Hopf fibration”), just by multiplying every point by $\bf{q}$}. However multiplying on the left (${\bf{x}}\to{\bf{q}}{\bf{x}}$) slides along rails in right-handed [sic] coils and multiplying on the right (${\bf{x}}\to{\bf{x}}{\bf{q}}$) slides along left-handed coils (as measured against a right-handed ordering of $\bf{i}$}, $\bf{j}$}, and $\bf{k}$}). We can think of these coiling around two special “straight” rails, the one passing through $\bf{1}$} and $\bf{q}$}, and one opposite one, the furthest one away (these are special only in our coordinates though and geometrically all the opposite pairs of rails are equal). As $S^{3}$ slides one full circuit along a special rail, it makes a full revolution around it (the same thing as making a full circuit along its opposite rail).
Figure 7: Models of symmetry types (left) $\pm[C_{2}\times C_{11}]$, (middle) $\pm[T\times C_{6}]$, six copies of a 16-cell slid along yellow rails, and (right) $O^{*}$ with edges corresponding to four-fold rotations of the cube.
Or we may specify a fibration of $S^{3}$ into great circle rails by giving some unit $\bf{q}$} on the rail through $\bf{1}$}. We can choose that $\bf{q}$} to be imaginary, and therefore on the unit great sphere midway between $\bf{1}$} and $\bf{1}$}. Multiplying by ${\bf{r}}:=(\cos\theta){\bf{1}}+(\sin\theta){\bf{q}}$ moves the hypersphere $\frac{\theta}{2\pi}$th the way along this rail, but twisting with a handedness: Right multiplication ${\bf{x}}\to{\bf{x}}{\bf{r}}$ rotates $S^{3}$ by $\theta$ clockwise, as seen from $\bf{r}$} looking to $\bf{1}$}; and left-multiplication ${\bf{x}}\to{\bf{r}}{\bf{x}}$ rotates $S^{3}$ by $\theta$ counter-clockwise. Therefore taking ${\bf{x}}\to{{\bf{r}}}^{-1}{\bf{x}}{\bf{r}}$ slides $S^{3}$ along, then back, fixing $\bf{1}$} but rotating $S^{3}$ about the $\bf{1}$}-to-$\bf{r}$} axis clockwise by $2\theta=2\arccos({\text{re}}({\bf{r}}))$. We can thusly encode three-dimensional rotations by unit quaternions $\bf{r}$}. The correspondence is two-to-one, though, because ${{\bf{r}}}^{-1}{\bf{x}}{\bf{r}}={(-{\bf{r}})}^{-1}{\bf{x}}(-{\bf{r}})$ and so $\bf{r}$} and $\bf{r}$} give the same rotation of $\bf{1}$}. Each 3-d rotational symmetry group $G$ lifts to a symmetrical set points $G^{*}$ in $S^{3}$ (in binary, opposite pairs). Multiplying on the right (say) by any of these points $\bf{q}$} rotates $S^{3}$ taking $\bf{1}$} to $\bf{q}$} and preserving $G^{*}$ as a whole — $G^{*}$ is a fixed-point free symmetry group of $G^{*}$ itself. V. Hart et al. cleverly show these “binary” rotation groups of the platonic solids with monkeys [7].
As an example, orient a tetrahedron so that its vertices are $(1,1,1)$ and permutations of (1,1,1)(-1,-1,1), (that is, every other vertex of the unit cube) and let $T$ be its rotational symmetry group. It happens that the twenty-four points $V_{24}$ are the binary tetrahedral group $T^{*}$: The four pairs of $V_{8}$ correspond to the identity and the three 2-fold rotations, and the eight pairs of $V_{16}$ correspond to the eight 3-fold rotations. In fact, edges correspond to rotations: If ${\bf{p}},{\bf{r}}\in V_{24}$ are the ends of an edge then ${\bf{e}}:=({\bf{p}}{\bf{r}}^{-1})$ corresponds to a three-fold rotation (that is ${\bf{e}}^{3}=-{\bf{1}}$), and vice versa. Moreover rotations that have the same axis correspond to edges lying on the same great circle paths. Figure 2 (d) shows this on 4DLO. At right in Figure 4 we color cycles of 24-cell edges by which three-fold rotation axis they correspond to, showing symmetry $T^{*}$.
We can map the 24-cell to its dual, taking $V_{24}$ to $V_{24}^{\prime}$ by multiplying (right or left) by $\frac{1}{\sqrt{2}}({\bf{1}}+{\bf{i}})$, or by any other point in $V_{24}^{\prime}$. The twelve pairs of points in $V_{24}^{\prime}$ each correspond to a quarter turn about a coordinate axis, exactly the remaining rotational symmetries $O$ of a cube, and $O^{*}=V_{24}\cup V_{24}$. Connecting these vertices by edges corresponding to four-fold rotations we have the model at lower left in Figure 5, or if we color these by rotation axis, we have the one at right in Figure 7. Directing these edges gives symmetry $O^{*}$, but these models have more symmetry, though, because we can pivot around points:
By multiplying on both left and right, we can fix points, and therefore have symmetry about them. By writing $\pm[A\times B]$ Conway and Smith mean the group of actions of the ${\bf{x}}\to{\bf{a}}{\bf{x}}{\bf{b}}$ where $\bf{a}$} and $\bf{b}$} map into 3d rotation groups $A$ and $B$. Figure 7 shows models of symmetry $\pm[C_{2}\times C_{11}]$ and $\pm[T\times C_{6}]$, where $C_{n}$ is the cyclic group of order $n$. That notation soon has many amendments — we rely on the detailed geometric discussion in  [11], especially their Sections 8.4 and 8.6 on the tesseract and 24-cell. We’ll give Conway and Smith’s notation, referring the reader to the concordances and extensions in [3, 4, 11].
As there is cubic symmetry around each point in $V_{24}\cup V_{24}^{\prime}$, this compound of two 24-cells has rotational symmetry $\pm[O\times O]$. Applying any ${\bf{x}}\to{\bf{r}}{\bf{x}}{\bf{p}}=({\bf{r}}{\bf{p}})({\bf{p}}^{-1}{\bf{x}}{\bf{p}})$ rotates $\bf{1}$} about the $\bf{p}$} axis (a rotation fixing $\bf{1}$} that is a symmetry in $O$) and then by left multiplication moves everything taking $\bf{1}$} to ${\bf{r}}{\bf{p}}^{-1}$, another point of $V_{24}\cup V_{24}^{\prime}$, also in $O^{*}$. The model at bottom left of Figure 5 also has vertices $V_{24}\cup V_{24}^{\prime}$, and has this same symmetry, but it is not two dual 24-cells.
Allowing mirror reflections by taking symmetries of the form ${\bf{x}}\to{\bf{p}}\overline{{\bf{x}}}{\bf{r}}$ where $\overline{{\bf{x}}}=2{\text{re}}({\bf{x}})-{\bf{x}}$, we extend these to the full symmetry of this dual pair, named $\pm[O\times O]\cdot 2$. Henceforth we’ll name only the rotational symmetries.
With half of the symmetry of a dual pair, 4DLO shows the full symmetry of a single 24-cell in Figure 2(a), denoted $\pm\frac{1}{2}[O\times O]$. We’ve seen that the vertices $V_{24}$ of a vertex-down 24-cell sort into $V_{8}$, $V_{16}^{+}$ and $V_{16}^{-}$. Any point corresponding to 3-fold rotation, so it happens, is a symmetry that permutes these sets. For example, multiplying by $\omega:=\frac{1}{2}({\bf{i}}+{\bf{j}}+{\bf{k}}+{\bf{1}})$ moves ${\bf{1}}\in V_{8}$ to $\omega\in V_{16}^{+}$ to $\omega^{2}\in V_{16}^{-}$ to ${\bf{1}}\in V_{8}$, etc. We can put a direction on the edges of a 24-cell in this way, shown in Figures 9 (left) and 2 (b) with type (we think) $\pm[T\times T]$.
Because $V_{8}$ is the vertices of a 16-cell, its images $V_{16}^{+}$ and $V_{16}^{-}$ are too, explaining the symmetry in the compound of three 16-cells, shown in Figures 5 and 6, with vertices $V_{24}$. Equivalently, the 16-cell, as well as its dual, the tesseract, has a third of the symmetry of a 24-cell, which Conway and Smith list as $\pm\frac{1}{6}[O\times O]$. Thus too do we find the compound of three tesseracts sharing vertices with a (dual) 24-cell, shown cell-down in Figure 4, or vertex down in 4DLO at left in Figure 1. There is a thus a compound of two compounds of three tesseracts, with the full symmetry $\pm[O\times O]$.
[…]
Figure 9: (left) From the interactive javascript model at chaimgoodmanstrauss.com/4DLO
(right) A prototype lighted 16-cell (Fayetteville, Ark., 2019)
Inspired by Leo Villareal’s Buckyball and Hive: Bleeker Street, we turned to moving lights to illustrate these symmetries, with a first project shown at right in Figure 9. The 4DLO was much more ambitious, consisting of about 300 feet of 1 inch outer diameter polyethylene tubing (with its beautiful capacity for radiating soft light), and 14,000 WS2811b LED lights diffused by strips of packing foam. These are controlled very simply – each light on each strand receives a string of bytes, taking three for its R, G, and B and passing the rest along. The 96 tubes of the model were organized into four quadrants, each with seven separate strands, powered at each node they passed through. A Teensy 4.1 has more than enough processing power and pins for this task, especially when decked out with a full suite of products from its producer PJRC Electronic Products: Four Octos, one for each quarter for the data lines, their Audio Shield for sound, and extra on-board memory. Each individual LED was associated with a particular location in 4D space, measured along the arc of an edge of the 24-cell; at each moment of the program, an edge would be associated with a coloring function, and these would be assigned by using whatever symmetry we wished to highlight. The javascript program at https://chaimgoodmanstrauss.com/4DLO/ allowed us to prototype the geometry and symmetry — you can also rotate the 24-cell into different positions. We used a bespoke scripting language for the sequencing, which we developed making extensive use of Claude.ai (along the way, learning much about the potential and limitations of today’s LLMs). All of our code is at the Github repository github.com/chaimgoodmanstrauss/4DLO.git. Needless to say, we made many mistakes, and learned a lot, much too much to fit here. (We printed four sets of nodes before getting it right.) Give us a call if you would like to know more.
Acknowledgements
This work has been supported by the Gathering for Gardner and the National Museum of Mathematics; and by the Institut Henri Poincaré (UAR 839 CNRS-Sorbonne Université, Labex Carmin ANR-10-LABX-59-01’); we are especially grateful to unnamed donors, to Cindy Lawrence, and to the MoMath staff for helping bringing 4DLO to life. The first author expresses deep appreciation for the second author’s ability to make these things real.